# Improving MLLMs in Embodied Exploration and Question Answering with Human-Inspired Memory Modeling

### v1

@@ IV-C Ablation Study
Table III presents ablation results on the A-EQA benchmark. When episodic memory is removed, we observe a clear degradation in exploration efficiency. Specifically, LLM-Match drops from 41.3% to 39.1% under the Qwen backbone, and from 65.6% to 60.9% under the GPT-4o backbone. This performance decrease indicates that our episodic memory provides effective guidance for exploration by enabling the agent to reuse relevant prior experiences. In contrast, removing semantic memory results in a more decline in answer accuracy than corresponding reduction in SPL. This highlights the critical role of semantic memory in preserving transferable and high-level rules and supporting embodied reasoning during question answering. Overall, these ablation results demonstrate that the two memory modules offer complementary benefits: episodic memory primarily improves exploration efficiency, while semantic memory contributes more substantially to accurate and reasoning-intensive question answering.
@@ V Conclusion
We presented a non-parametric memory framework for MLLM-based embodied agents that explicitly disentangles episodic and semantic memory to address long-horizon exploration and reasoning. By adopting a visual reasoning paradigm, our approach enables agents to recall and verify prior episodic experiences without relying on rigid geometric fusion, thereby supporting robust cross-episode reuse of observations and improving exploration efficiency. In addition, resuable and structed extracted rules from experiences tackle complex decision making in embodied tasks. Comprehensive evaluation on A-EQA and GOAT-Bench demonstrates the effectiveness of our memory modeling in improving both exploration efficiency and answering accuracy. We identify our limitations points to promising future directions: 1) Differences in scale, map and visual may introduce new challenges for episodic memory in outdoor environments. 2) Parametrically integrating semantic memory into a embodied foundation model for joint optimization of memory recalling.

### v2

@@ IV-C Ablation Study
Table III presents ablation results on the A-EQA benchmark. When episodic memory is removed, we observe a clear degradation in exploration efficiency. Specifically, LLM-Match drops from 41.3% to 39.1% under the Qwen backbone, and from 65.6% to 60.9% under the GPT-4o backbone. This performance decrease indicates that our episodic memory provides effective guidance for exploration by enabling the agent to reuse relevant prior experiences. In contrast, removing semantic memory results in a more decline in answer accuracy than corresponding reduction in SPL. This highlights the critical role of semantic memory in preserving transferable knowledge and supporting embodied reasoning during question answering. Overall, these ablation results demonstrate that the two memory modules offer complementary benefits: episodic memory primarily improves exploration efficiency, while semantic memory contributes more substantially to accuracy of reasoning-intensive question answering.
@@ V Conclusion
We presented HIMM, a novel memory framework for MLLM-based embodied agents that explicitly model episodic and semantic memory to address long-horizon exploration and reasoning. By constructing a visual reasoning paradigm, our approach enables agents to recall and verify past episodic experiences without relying on rigid geometric fusion, thereby supporting robust cross-episode reuse of observations and improving exploration efficiency. In addition, resuable and structed rules as semantic memory distilled from experiences tackle complex reasoning in embodied tasks. Comprehensive evaluation on A-EQA and GOAT-Bench demonstrates the effectiveness of our memory modeling in improving both exploration efficiency and answering accuracy.